# OmnimatteZero: Training-free Real-time Omnimatte with Pre-trained Video Diffusion Models

### v1

@@ 5.3 Layer composition
The objective here is to take the extracted foreground layers and seamlessly composite them onto new background layers. Figure 7 presents the results of OmnimatteZero, highlighting its ability to maintain object integrity, shadows, and reflections while enabling smooth re-composition across different scenes. Notably, it demonstrates how the model adaptively integrates the inserted object into the environment, such as enhancing the cat’s reflection in the clear water stain or adjusting the man’s appearance to match the night lighting.
6 Summary
We present the first real-time, training-free generative approach for layered video decomposition. Our method extracts clean background layers and foreground objects while preserving effects like shadows and reflections, enabling seamless re-composition without dataset-specific optimization. Evaluations demonstrate that OmnimatteZero outperforms all training based methods as well as optimization based methods in terms of background reconstruction and running time. However, some challenges persist, as the output can be sensitive to the quality of VAE video encodings, potentially leading to slight differences from the original video (see “running dog” example in supp). With advancements in video diffusion models, we anticipate that these issues will be addressed in the near future.

### v2

@@ 6.3. Layer composition
The objective here is to take the extracted foreground layers and seamlessly composite them onto new background layers. Figures 8 and 11 presents the results of OmnimatteZero, highlighting its ability to maintain object integrity, shadows, and reflections while enabling smooth re-composition across different scenes. Notably, it demonstrates how the model adaptively integrates the inserted object into the environment, such as enhancing the cat’s reflection in the clear water stain or adjusting the man’s appearance to match the night lighting.
6.4. Acknowledgement
We thank Yoni Kasten for his insightful input and valuable contributions to this work.
Appendix A Implementation Details:
We apply OmnimatteZero using two video diffusion models: LTXVideo (HaCohen et al., 2024) v0.9.1 and Wan2.1 (video, 2025), both running with $T_{\text{steps}}=30$ denoising steps. The guidance scale was set to 0 (i.e. no prompt) as we found no major effect of the prompt on the process, and all result videos were generated using the same random seed.
Quantitative and qualitative results of baselines were obtained from the respective authors. All qualitative results in this paper are based on LTXVideo, as it shows minimal visual differences from Wan2.1. Additionally, due to space constraints, each figure displays a single frame per video for qualitative results. Full videos are available in the supplementary material.
Latent masking: Applying a given mask video directly in latent space is challenging due to the unclear mapping from pixel space to the latent space. We start by taking the object masks (per-frame) and computing a binary video, where the frames are binary images, and pass this video as an RGB input (with duplicated channels) through the VAE encoder as well. This process produces a latent representation where high values indicate the masked object, enabling its spatio-temporal footprint identification via simple thresholding.
Appendix B Ablation Study
To assess the individual contributions of our key architectural components, we perform a series of ablation studies on the Movie benchmark using the LTXVideo dataset (HaCohen et al., 2024). Quantitative results are reported in terms of PSNR.
B.1. Effect of Temporal and Spatial Attention Guidance
We first analyze the impact of temporal and spatial attention mechanisms on reconstruction quality:
Temporal Spatial PSNR
18.5
28.3
23.2
35.11 (Ours)
Table 2. Effect of temporal and spatial attention guidance. Temporal guidance contributes most significantly, with the highest quality achieved when both cues are combined.
Results indicate that temporal attention is the primary contributor to reconstruction fidelity, with a PSNR improvement of nearly 10 points over the baseline. Spatial attention offers complementary benefits. When integrated, the two mechanisms produce a substantial synergistic effect, resulting in the highest overall performance.
B.2. Influence of Point Sampling in Track-Any-Point
Next, we evaluate the effect of varying the sampling density of points from the object mask in the Track-Any-Point module:
Percent of Object Pixels PSNR
20% 22.6
40% 28.1
60% 35.11
80% 35.11
100% 35.12
Table 3. Effect of point sampling density. Using 60% of object pixels is sufficient to match the maximum reconstruction quality.
We observe that increasing the proportion of sampled object pixels leads to improved reconstruction quality, with performance saturating at approximately 60% coverage. This suggests that high-fidelity results can be attained with moderately dense sampling, highlighting the efficiency of our approach in establishing accurate correspondences.
B.3. Impact of Effect-Aware Masking
We investigate the influence of explicitly incorporating associated object-induced visual effects—such as shadows and reflections—into the masking process:
Effect Masking PSNR
Without 29.16
With (Ours) 35.11
Table 4. Effect of including associated visual effects (e.g., shadows, reflections) in the mask. These effects are essential for clean background reconstruction.
The inclusion of effect-aware masks yields a substantial improvement of +6.0 PSNR, highlighting the critical role of accurately capturing the extended influence of foreground objects on their surrounding environment. This result underscores the importance of modeling object footprints for realistic scene reconstruction.
[…]
[Description]Description
Figure 9. Qualitative comparison with state-of-the-art video inpainting methods. OmnimatteZero produces clean, temporally consistent background reconstructions across a diverse range of scenes, including dynamic motion, shadows, reflections, and fine textures. Compared to other methods, it avoids common artifacts such as ghosting, blur, and shadow remnants (highlighted in red boxes), successfully filling in complex backgrounds across varying temporal and spatial contexts. Gray frames indicate failure cases where the inpainting method returned an error. See the supplementary material for full videos.
[…]
[Description]Description
Figure 10. Qualitative comparison with state-of-the-art video inpainting methods. OmnimatteZero achieves clean, temporally consistent background reconstructions across diverse scenes, preserving fine textures, shadows, and reflections. Compared to prior methods, it avoids common artifacts such as ghosting, blur, and residual shadows (highlighted in red). Gray frames indicate failure cases where the method returned an error. See supplementary material for full video results.
[…]
[Description]Description
Figure 11. Qualitative Comparison: Layered Composition. From left to right: the original video frame, the extracted foreground (including shadows and reflections), the target background, and our final composite. Across four diverse examples (a swan, a cat, a bicyclist, and a dog), our method preserves accurate shadows, reflections, and object–scene coherence, yielding visually seamless results. See the supplementary material for full videos.